# Wireless Communication for Low-Altitude Economy with UAV Swarm Enabled Two-Level Movable Antenna System

### v1

@@ Appendix B Proof of Lemma 1
With ${g_{\_}{k,i,l}}$ given in (32), its lower and upper bounds can be constructed based on the second-order Taylor expansion [47, 24, 48]. Specifically, the gradient of ${g_{\_}{k,i,l}}$ over ${{{\bf{q}}_{\_}l}}$ is $\nabla{g_{\_}{k,i,l}}={\left[{\frac{{\partial{g_{\_}{k,i,l}}}}{{\partial{x_{\_}l}}},\frac{{\partial{g_{\_}{k,i,l}}}}{{\partial{y_{\_}l}}},\frac{{\partial{g_{\_}{k,i,l}}}}{{\partial{z_{\_}l}}}}\right]^{T}}$, where
$\frac{{\partial{g_{\_}{k,i,l}}}}{{\partial{x_{\_}l}}}=-\frac{{2\pi}}{\lambda}\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|}{\Phi_{\_}i}$ (56)
×$\ \ \ \ \ \ \ \ \ \ \times\sin\left({\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)-\angle{V_{\_}{k,l,l^{\prime}}}}\right),$
$\frac{{\partial{g_{\_}{k,i,l}}}}{{\partial{y_{\_}l}}}=-\frac{{2\pi}}{\lambda}\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|}{\Psi_{\_}i}$ (57)
×$\ \ \ \ \ \ \ \ \ \ \times\sin\left({\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)-\angle{V_{\_}{k,l,l^{\prime}}}}\right),$
$\frac{{\partial{g_{\_}{k,i,l}}}}{{\partial{z_{\_}l}}}=-\frac{{2\pi}}{\lambda}\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|}{\Theta_{\_}i}$ (58)
×$\ \ \ \ \ \ \ \ \ \ \times\sin\left({\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)-\angle{V_{\_}{k,l,l^{\prime}}}}\right),$
Besides, the Hessian matrix of ${{g_{\_}{k,i,l}}}$ over ${{{\bf{q}}_{\_}l}}$ is
@@
$\frac{{{\partial^{2}}{g_{\_}{k,i,l}}}}{{\partial{x_{\_}l}\partial{x_{\_}l}}}=-{\left({\frac{{2\pi}}{\lambda}}\right)^{2}}\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|}\Phi_{\_}i^{2}$ (60)
×$\times\cos\left({\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)-\angle{V_{\_}{k,l,l^{\prime}}}}\right),$
@@
$\frac{{{\partial^{2}}{g_{\_}{k,i,l}}}}{{\partial{y_{\_}l}\partial{y_{\_}l}}}=-{\left({\frac{{2\pi}}{\lambda}}\right)^{2}}\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|}\Psi_{\_}i^{2}$ (61)
×$\times\cos\left({\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)-\angle{V_{\_}{k,l,l^{\prime}}}}\right),$
$\frac{{{\partial^{2}}{g_{\_}{k,i,l}}}}{{\partial{z_{\_}l}\partial{z_{\_}l}}}=-{\left({\frac{{2\pi}}{\lambda}}\right)^{2}}\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|}\Theta_{\_}i^{2}$ (62)
×$\times\cos\left({\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)-\angle{V_{\_}{k,l,l^{\prime}}}}\right),$
@@
$\frac{{{\partial^{2}}{g_{\_}{k,i,l}}}}{{\partial{x_{\_}l}\partial{y_{\_}l}}}$ $=\frac{{{\partial^{2}}{g_{\_}{k,i,l}}}}{{\partial{y_{\_}l}\partial{x_{\_}l}}}=-{\left({\frac{{2\pi}}{\lambda}}\right)^{2}}\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|{\Phi_{\_}i}{\Psi_{\_}i}}$ (63)
×$\ \ \ \ \ \times\cos\left({\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)-\angle{V_{\_}{k,l,l^{\prime}}}}\right),$
@@
$\frac{{{\partial^{2}}{g_{\_}{k,i,l}}}}{{\partial{x_{\_}l}\partial{z_{\_}l}}}$ $=\frac{{{\partial^{2}}{g_{\_}{k,i,l}}}}{{\partial{z_{\_}l}\partial{x_{\_}l}}}=-{\left({\frac{{2\pi}}{\lambda}}\right)^{2}}\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|{\Phi_{\_}i}{\Theta_{\_}i}}$ (64)
×$\ \ \ \ \ \times\cos\left({\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)-\angle{V_{\_}{k,l,l^{\prime}}}}\right),$
@@
$\frac{{{\partial^{2}}{g_{\_}{k,i,l}}}}{{\partial{y_{\_}l}\partial{z_{\_}l}}}$ $=\frac{{{\partial^{2}}{g_{\_}{k,i,l}}}}{{\partial{z_{\_}l}\partial{y_{\_}l}}}=-{\left({\frac{{2\pi}}{\lambda}}\right)^{2}}\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|{\Psi_{\_}i}{\Theta_{\_}i}}$ (65)
×$\ \ \ \ \ \times\cos\left({\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right)-\angle{V_{\_}{k,l,l^{\prime}}}}\right).$
Moreover, with (59), it follows that
$\left\|{{\nabla^{2}}{g_{\_}{k,i,l}}}\right\|_{\_}2^{2}\leq\left\|{{\nabla^{2}}{g_{\_}{k,i,l}}}\right\|_{\_}F^{2}$ (66)
$\ \ \ ={\left({\frac{{{\partial^{2}}{g_{\_}{k,i,l}}}}{{\partial{x_{\_}l}\partial{x_{\_}l}}}}\right)^{2}}+\cdots+{\left({\frac{{{\partial^{2}}{g_{\_}{k,i,l}}}}{{\partial{z_{\_}l}\partial{z_{\_}l}}}}\right)^{2}}$
@@
$\ \ \ \leq{\left({\frac{{4{\pi^{2}}}}{{{\lambda^{2}}}}\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|}}\right)^{2}}\left({\Phi_{\_}i^{4}+\Phi_{\_}i^{2}\Psi_{\_}i^{2}+}\right.$
$\ \ \ \ \ \ \left.{\Phi_{\_}i^{2}\Theta_{\_}i^{2}+\Psi_{\_}i^{2}\Phi_{\_}i^{2}+\Psi_{\_}i^{4}+\Psi_{\_}i^{2}\Theta_{\_}i^{2}+\Theta_{\_}i^{2}\Phi_{\_}i^{2}+\Theta_{\_}i^{2}\Psi_{\_}i^{2}+\Theta_{\_}i^{4}}\right)$
$\ \ \ ={\left({\frac{{4{\pi^{2}}}}{{{\lambda^{2}}}}\sum\limits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{\_}i}}\right|}^{2}}\left|{{V_{\_}{k,l,l^{\prime}}}}\right|}}\right)^{2}}.$
@@
With ${\nabla^{2}}{g_{\_}{k,i,l}}\preceq{\left\|{{\nabla^{2}}{g_{\_}{k,i,l}}}\right\|_{\_}2}{\bf{I}}$, and by choosing ${\delta_{\_}{k,i,l}}=\frac{{8{\pi^{2}}}}{{{\lambda^{2}}}}{\left|{{\alpha_{\_}i}}\right|^{2}}\sum\nolimits_{\_}{l^{\prime}=1,l^{\prime}\neq l}^{L}{\left|{{V_{\_}{k,l,l^{\prime}}}}\right|}$, we have ${\nabla^{2}}{g_{\_}{k,i,l}}\preceq{\delta_{\_}{k,i,l}}{\bf{I}}$. Thus, the lower and upper bounds of ${{g_{\_}{k,i,l}}}$ in (33) and (34) can be obtained with the Taylor’s theorem [47]. The proof of Lemma 1 is thus completed.

### v2

@@ Appendix B Proof of Lemma 1
With $g_{k,i,l}$} given in (31), its lower and upper bounds can be constructed based on the second-order Taylor expansion [48, 24, 49]. Specifically, the gradient of $g_{k,i,l}$} over ${\bf{q}}_{l}$}} is $\nabla{g_{k,i,l}}={\left[{\frac{{\partial{g_{k,i,l}}}}{{\partial{x_{l}}}},\frac{{\partial{g_{k,i,l}}}}{{\partial{y_{l}}}},\frac{{\partial{g_{k,i,l}}}}{{\partial{z_{l}}}}}\right]^{T}}$, where
$\frac{{\partial{g_{k,i,l}}}}{{\partial{x_{l}}}}=-\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|\frac{{\partial{r_{i,l}}}}{{\partial{x_{l}}}}}$ (55)
$\ \ \ \ \ \ \ \ \ \ \times{\rm{sin}}\left({\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)+\angle{V_{k,l,l^{\prime}}}}\right),$
$\frac{{\partial{g_{k,i,l}}}}{{\partial{y_{l}}}}=-\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|\frac{{\partial{r_{i,l}}}}{{\partial{y_{l}}}}}$ (56)
$\ \ \ \ \ \ \ \ \ \ \times{\rm{sin}}\left({\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)+\angle{V_{k,l,l^{\prime}}}}\right),$
$\frac{{\partial{g_{k,i,l}}}}{{\partial{z_{l}}}}=-\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|\frac{{\partial{r_{i,l}}}}{{\partial{z_{l}}}}}$ (57)
$\ \ \ \ \ \ \ \ \ \ \times{\rm{sin}}\left({\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)+\angle{V_{k,l,l^{\prime}}}}\right),$
where $\frac{{\partial{r_{i,l}}}}{{\partial{x_{l}}}}={\Phi_{i}}+\frac{{{x_{l}}-\left({{x_{l}}{\Phi_{i}}+{y_{l}}{\Psi_{i}}+{z_{l}}{\Theta_{i}}}\right){\Phi_{i}}}}{{{r_{i}}}}$, $\frac{{\partial{r_{i,l}}}}{{\partial{y_{l}}}}={\Psi_{i}}+\frac{{{y_{l}}-\left({{x_{l}}{\Phi_{i}}+{y_{l}}{\Psi_{i}}+{z_{l}}{\Theta_{i}}}\right){\Psi_{i}}}}{{{r_{i}}}}$, and $\frac{{\partial{r_{i,l}}}}{{\partial{z_{l}}}}={\Theta_{i}}+\frac{{{z_{l}}-\left({{x_{l}}{\Phi_{i}}+{y_{l}}{\Psi_{i}}+{z_{l}}{\Theta_{i}}}\right){\Theta_{i}}}}{{{r_{i}}}}$, respectively.
Besides, the Hessian matrix of $g_{k,i,l}$}} over ${\bf{q}}_{l}$}} is
@@
$\frac{{{\partial^{2}}{g_{k,i,l}}}}{{\partial{x_{l}}\partial{x_{l}}}}=-\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|\times}$ (59)
$\ \ \left[{\frac{{1-\Phi_{i}^{2}}}{{{r_{i}}}}{\rm{sin}}\left({\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)+\angle{V_{k,l,l^{\prime}}}}\right)+}\right$.
$\ \ \ \left.{\frac{{2\pi}}{\lambda}{{\left({\frac{{\partial{r_{i,l}}}}{{\partial{x_{l}}}}}\right)}^{2}}\cos\left({\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)+\angle{V_{k,l,l^{\prime}}}}\right)}\right],$
@@
$\frac{{{\partial^{2}}{g_{k,i,l}}}}{{\partial{x_{l}}\partial{y_{l}}}}=-\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|\times}$ (60)
$\ \ \left[{\frac{{-{\Phi_{i}}{\Psi_{i}}}}{{{r_{i}}}}{\rm{sin}}\left({\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)+\angle{V_{k,l,l^{\prime}}}}\right)+}\right$.
$\ \ \ \left.{\frac{{2\pi}}{\lambda}\frac{{\partial{r_{i,l}}}}{{\partial{x_{l}}}}\frac{{\partial{r_{i,l}}}}{{\partial{y_{l}}}}\cos\left({\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)+\angle{V_{k,l,l^{\prime}}}}\right)}\right],$
@@
$\frac{{{\partial^{2}}{g_{k,i,l}}}}{{\partial{x_{l}}\partial{z_{l}}}}=-\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|\times}$ (61)
$\ \ \left[{\frac{{-{\Phi_{i}}{\Theta_{i}}}}{{{r_{i}}}}{\rm{sin}}\left({\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)+\angle{V_{k,l,l^{\prime}}}}\right)+}\right$.
$\ \ \ \left.{\frac{{2\pi}}{\lambda}\frac{{\partial{r_{i,l}}}}{{\partial{x_{l}}}}\frac{{\partial{r_{i,l}}}}{{\partial{z_{l}}}}\cos\left({\frac{{2\pi}}{\lambda}\left({{r_{i,l}}-{r_{i,l^{\prime}}}}\right)+\angle{V_{k,l,l^{\prime}}}}\right)}\right],$
and other elements can be similarly obtained based on (55)-(57), which are omitted for brevity.
Moreover, with (58), we have (62), shown at the top of the next page.
@@
$\left\|{{\nabla^{2}}{g_{k,i,l}}}\right\|_{2}^{2}\leq\left\|{{\nabla^{2}}{g_{k,i,l}}}\right\|_{F}^{2}\;={\left({\frac{{{\partial^{2}}{g_{k,i,l}}}}{{\partial{x_{l}}\partial{x_{l}}}}}\right)^{2}}+\cdots+{\left({\frac{{{\partial^{2}}{g_{k,i,l}}}}{{\partial{z_{l}}\partial{z_{l}}}}}\right)^{2}}\leq{\left({\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}}\right)^{2}}\times$ (62)
$\left[{{{\left({\frac{{1-\Phi_{i}^{2}}}{{{r_{i}}}}}\right)}^{2}}+{{\left({\frac{{2\pi}}{\lambda}{{\left({\frac{{\partial{r_{i,l}}}}{{\partial{x_{l}}}}}\right)}^{2}}}\right)}^{2}}+{{\left({\frac{{{\Phi_{i}}{\Psi_{i}}}}{{{r_{i}}}}}\right)}^{2}}+{{\left({\frac{{2\pi}}{\lambda}\frac{{\partial{r_{i,l}}}}{{\partial{x_{l}}}}\frac{{\partial{r_{i,l}}}}{{\partial{y_{l}}}}}\right)}^{2}}+{{\left({\frac{{{\Phi_{i}}{\Theta_{i}}}}{{{r_{i}}}}}\right)}^{2}}+{{\left({\frac{{2\pi}}{\lambda}\frac{{\partial{r_{i,l}}}}{{\partial{x_{l}}}}\frac{{\partial{r_{i,l}}}}{{\partial{z_{l}}}}}\right)}^{2}}+}\right$.
${\left({\frac{{{\Psi_{i}}{\Phi_{i}}}}{{{r_{i}}}}}\right)^{2}}+{\left({\frac{{2\pi}}{\lambda}\frac{{\partial{r_{i,l}}}}{{\partial{y_{l}}}}\frac{{\partial{r_{i,l}}}}{{\partial{x_{l}}}}}\right)^{2}}+{\left({\frac{{1-\Psi_{i}^{2}}}{{{r_{i}}}}}\right)^{2}}+{\left({\frac{{2\pi}}{\lambda}{{\left({\frac{{\partial{r_{i,l}}}}{{\partial{y_{l}}}}}\right)}^{2}}}\right)^{2}}+{\left({\frac{{{\Psi_{i}}{\Theta_{i}}}}{{{r_{i}}}}}\right)^{2}}+{\left({\frac{{2\pi}}{\lambda}\frac{{\partial{r_{i,l}}}}{{\partial{y_{l}}}}\frac{{\partial{r_{i,l}}}}{{\partial{z_{l}}}}}\right)^{2}}+$
$\left.{{{\left({\frac{{{\Theta_{i}}{\Phi_{i}}}}{{{r_{i}}}}}\right)}^{2}}+{{\left({\frac{{2\pi}}{\lambda}\frac{{\partial{r_{i,l}}}}{{\partial{z_{l}}}}\frac{{\partial{r_{i,l}}}}{{\partial{x_{l}}}}}\right)}^{2}}+{{\left({\frac{{{\Theta_{i}}{\Psi_{i}}}}{{{r_{i}}}}}\right)}^{2}}+{{\left({\frac{{2\pi}}{\lambda}\frac{{\partial{r_{i,l}}}}{{\partial{z_{l}}}}\frac{{\partial{r_{i,l}}}}{{\partial{y_{l}}}}}\right)}^{2}}+{{\left({\frac{{1-\Theta_{i}^{2}}}{{{r_{i}}}}}\right)}^{2}}+{{\left({\frac{{2\pi}}{\lambda}{{\left({\frac{{\partial{r_{i,l}}}}{{\partial{z_{l}}}}}\right)}^{2}}}\right)}^{2}}}\right]$
@@
$={\left({\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}}\right)^{2}}\left[{2+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{x_{l}^{2}+y_{l}^{2}+z_{l}^{2}-{{\left({{x_{l}}{\Phi_{i}}+{y_{l}}{\Psi_{i}}+{z_{l}}{\Theta_{i}}}\right)}^{2}}}}{{r_{i}^{2}}}}\right)}^{2}}}\right].$
Let ${C_{\max}}\triangleq\mathop{\max}\limits_{{{\bf{q}}_{l}}\in{\cal C}}x_{l}^{2}+y_{l}^{2}+z_{l}^{2}$, it follows that
@@
$\left\|{{\nabla^{2}}{g_{k,i,l}}}\right\|_{F}^{2}$ $\leq{\left({\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}}\right)^{2}}\times$ (63)
$\ \ \ \ \ \left[{2+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{{C_{\max}}}}{{r_{i}^{2}}}}\right)}^{2}}}\right].$
@@
With ${\nabla^{2}}{g_{k,i,l}}\preceq{\left\|{{\nabla^{2}}{g_{k,i,l}}}\right\|_{2}}{\bf{I}}$, and by choosing ${\delta_{k,i,l}}=\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}\sqrt{2+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{{C_{\max}}}}{{r_{i}^{2}}}}\right)}^{2}}}$, we have ${\nabla^{2}}{g_{k,i,l}}\preceq{\delta_{k,i,l}}{\bf{I}}$. Thus, the lower and upper bounds of $g_{k,i,l}$}} in (32) and (33) can be obtained with the Taylor’s theorem [48]. The proof of Lemma 1 is thus completed.
Appendix C Proof of Proposition 1
Denote by $\{{{\bf{q}}_{l}^{\left(j\right)}}\}$ and $\{{{\bf{v}}_{k}^{\left(j\right)}}\}$ the corresponding optimization variables in the $j$-th iteration. Let $\rho_{{\rm{lb}}}^{{\rm{traj}}}({\{{{\bf{q}}_{l}^{\left(j\right)}}\},\{{{\bf{v}}_{k}^{\left(j\right)}}\}})$ denote the corresponding objective value of problem (38). In the $j$-th iteration, since the optimal solution to (24) is obtained for given $\{{\bf{q}}_{l}^{\left(j\right)}\}$ in step 3, it follows that
$\rho\left({\left\{{{\bf{q}}_{l}^{\left(j\right)}}\right\},\left\{{{\bf{v}}_{k}^{\left(j\right)}}\right\}}\right)\leq\rho\left({\left\{{{\bf{q}}_{l}^{\left(j\right)}}\right\},\left\{{{\bf{v}}_{k}^{\left({j+1}\right)}}\right\}}\right).$ (64)
Besides, regarding UAV $l$ in step 5, for given $\{{\bf{v}}_{k}^{\left({j+1}\right)}\}$ and $\{{\bf{q}}_{1}^{(j+1)},\cdots,{\bf{q}}_{l-1}^{(j+1)},{\bf{q}}_{l}^{(j)},\cdots,{\bf{q}}_{L}^{(j)}\}$, we have
$\rho\left({\left\{{{\bf{q}}_{1}^{(j+1)},\cdots,{\bf{q}}_{l-1}^{(j+1)},{\bf{q}}_{l}^{(j)},\cdots,{\bf{q}}_{L}^{(j)}}\right\},\left\{{{\bf{v}}_{k}^{\left({j+1}\right)}}\right\}}\right)$ (65)
$\mathop{=}\limits^{\left(a\right)}\rho_{{\rm{lb}}}^{{\rm{traj}}}\left({\left\{{{\bf{q}}_{1}^{(j+1)},\cdots,{\bf{q}}_{l-1}^{(j+1)},{\bf{q}}_{l}^{(j)},\cdots,{\bf{q}}_{L}^{(j)}}\right\},\left\{{{\bf{v}}_{k}^{\left({j+1}\right)}}\right\}}\right)$
$\mathop{\leq}\limits^{\left(b\right)}\rho_{{\rm{lb}}}^{{\rm{traj}}}\left({\left\{{{\bf{q}}_{1}^{(j+1)},\cdots,{\bf{q}}_{l-1}^{(j+1)},{\bf{q}}_{l}^{(j+1)},\cdots,{\bf{q}}_{L}^{(j)}}\right\},\left\{{{\bf{v}}_{k}^{\left({j+1}\right)}}\right\}}\right)$
$\mathop{\leq}\limits^{\left(c\right)}\rho\left({\left\{{{\bf{q}}_{1}^{(j+1)},\cdots,{\bf{q}}_{l-1}^{(j+1)},{\bf{q}}_{l}^{(j+1)},\cdots,{\bf{q}}_{L}^{(j)}}\right\},\left\{{{\bf{v}}_{k}^{\left({j+1}\right)}}\right\}}\right),$
where $\left(a\right)$ holds because the Taylor expansions are tight at the given local point ${\bf{q}}_{l}^{(j)}$; the inequality $\left(b\right)$ holds because problem (38) is solved via SCA, which ensures a non-decreasing objective value; the inequality $\left(c\right)$ holds because the objective value of (26) is lower-bounded by that of (38) at ${\bf{q}}_{l}^{(j+1)}$. Based on (64) and (65), we have
$\rho\left({\left\{{{\bf{q}}_{l}^{\left(j\right)}}\right\},\left\{{{\bf{v}}_{k}^{\left(j\right)}}\right\}}\right)\leq\rho\left({\left\{{{\bf{q}}_{l}^{\left({j+1}\right)}}\right\},\left\{{{\bf{v}}_{k}^{\left({j+1}\right)}}\right\}}\right).$ (66)
The result shows that the objective value of problem (12) is non-decreasing over iterations, and thus, Algorithm 1 is guaranteed to converge. This thus completes the proof of Proposition 1.